Shift report 20000629 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Jun 29 2000 - 08:01:35 EDT

  • Next message: http account: "Shift report 20000629 08:00-16:00"

    ------------------------------------------------------------
    Date: Thu Jun 29 08:01:35 2000, Shift: 00:00-08:00
    Supervisor: KH, with: AM
    Log book#2 pages 93 - 93
    
    Shift Summary:
    ------------------------------------------------------------
    No beam at all.
    There was a "quench" event during the last shift. 
    But starting at 2:00 monitor said quench event had
    been cleared and beam was expected by 3:30.  Never
    happened.
    TPC anode voltages turned down.
    No noticable detector problems
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Jun 29 2000 - 08:05:48 EDT